The probability that the event A occurs is Pr( A ) = 1/3 and the probability that the union event A∪B occurs is Pr( A∪B ) = 5/6. Answer the following questions.

1. If the events A and B are disjoint, what is the value of Pr(B)?

2. If the events A and B are independent, what is the value of Pr(B)?

(*Answer in the decimal form, not in the fractional form.)

Thank you so much!
Answer #1


P (A) = 1/3 ; PLAUB) = I P(AUB) = PIADA PCB) – PIANB) A and B are dos joint then Plans Now P(AUB) = PCAS+PCB) & = YztPIB) ) P